A student invested some money at an annual rate of 5%. If the annual income from the investment is $300, how much did he invest?
Found 2 solutions by rfer, jim_thompson5910:
Answer by rfer(16322)

300/0.05=$6000

Answer by jim_thompson5910(35256)

I = P*r*t

300 = P*0.05*1

300 = 0.05P

300/0.05 = P

6000 = P

P = 6000


He invested $6,000